Circular No. 7836

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002bf IN CGCG 266-031 P. Martin and W. D. Li, University of California at Berkeley, report the

LOTOSS (cf. IAUC 7514) discovery of an apparent supernova on unfiltered CCD images taken with the 0.8-m Katzman Automatic Imaging Telescope (KAIT) on Feb. 22.3 (mag ∼ 17.0) and 24.3 UT (mag ∼ 16.6). SN 2002bf is located at α = 10^{h}15^{m}42^{s}.31, δ = +55^{o}40′06′′.7 (equinox 2000.0), which is 0′′.6 east and 4′′.0 north of the nucleus of CGCG 266-031. A KAIT image taken on Feb. 3.3 showed nothing at this position (limiting mag ∼ 19.0).

SUPERNOVA 2002bg IN MCG +02-38-31 T. Puckett, Mountain Town, GA; and J. Newton, Chiefland, FL, re-

port the discovery of an apparent supernova (mag 16.9) on an unfiltered CCD frame (limiting mag 19.2) taken with the Puckett Observatory 0.35-m automated patrol telescope on Feb. 23.39 UT. SN 2002bg is located at α = 14^{h}59^{m}39^{s}.23, δ = +13^{o}12′44′′.5 (equinox 2000.0), which is 13′′.9 east and 9′′.9 south of the center of MCG +02-38-31. The new object was confirmed on CCD frames taken on Feb. 25.22 by Puckett with a 0.60-m reflector. The candidate is not present on Palomar Sky Survey images taken on 1993 Mar. 23, 1991 May 16 (limiting mag ∼ 21.0), and 1954 June 3 (limiting mag ∼ 19.7).

SUPERNOVA 2002ap IN M74 H. Yamaoka, Kyushu University, reports that he measured red mag

15.2 from an unfiltered prediscovery image of SN 2002ap obtained by W. Rogers (Endicott, NY) on Jan. 29.04 UT.

SUPERNOVA 2002ao IN UGC 9299 Yamaoka also notes that SN 2002ao has faded unusually rapidly, resem-

bling that of the type-Ic SN 1999cq (cf. IAUC 7825). Selected magnitudes (unfiltered CCD unless otherwise noted) reported to the VSNET collabolation: Jan. 28.175 UT, 14.8 (J.-M. Llapasset, Perpignan, France); 29.183, 14.8 (Llapasset); 31.7, ∼ 15.2 (K. Kinugasa and H. Kawakita, Gunma Astronomical Observatory; measured by Yamaoka); Feb. 13.207, 16.3 (Llapasset); 13.476, 16.4 (D. West, Mulvane, KS); 13.832, R = 16.58 (J. Watanabe, National Astronomical Observatory of Japan, Kiso 1.05-m Schmidt telescope + CCD; measured by Yamaoka); 13.847, R_{j} = 16.36 (H. Tsuda, Hokkaido, Japan; measured by Y. Sano, Nayoro, Japan).

2002 February 25  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green